In a small town in outback Australia the probability of a random resident supporting the Rams football team is 0.44, while the probability of a random resident supporting the Ewes football team is 0.39. These two sets are mutually exclusive. The probability of a random resident living on the south side of the river is 0.23, but this is independent of football affiliation (or lack thereof). What is the probability that a random resident will live on the north side of the river, but not follow a football team? (Answer to 2 decimal places)
Answer #1

P(South side) = 0.23

So,

P(North side) = 1 - 0.23 = 0.77

P(Follow a football team) = P(Supporting the Rams football team) + P(Supporting Ewes football team) = 0.44 + 0.39 = 0.83

So,

P(Not follow a football team) = 1 - 0.83 = 0.17

So,

P(North side & Not follow a football team) = 0.77 X 0.17 = 0.13

So,

Answer is:

0.13